write the equation of a line that is perpendicular to x + 4y=7 and through point(1,1)in slope intercept form
Answer by palanisamy(496)

The equation of the given line is x+4y = 7
The slope of the given line = -(coefficient of x)/(coefficient of y)=-1/4
The slope of a line perpendicular to the given line = 4
Therefore, the equation of a line that is perpendicular to x + 4y=7 and
through point(1,1) is y-1 = 4(x-1)
y-1 = 4x-4
y = 4x-3